Science Prep Night 20121202 UT

LBC

22:30 Started LBCs, DGH authorized, and connected

00:24 USpec and F972N20

USpec on Blue
003943 = 30000 counts
004022 = 26000 counts
004106 = 22000 counts
004154 = 18000 counts
004239 = 15500 counts

F972N20 on Red
003937 = 20000 counts
004022 = 17000 counts
004106 = 14000 counts
004151 = 12000 counts
004237 = 10000 counts

00:55 V-Bessel and V-Bessel; counts are fine, but some stars in the field. Assume dither will take care of this.

V-Bessel on Red
004822 = 20000 counts
004856 = 17000 counts
004940 = 14500 counts
005019 = 13000 counts
005058 = 10000 counts
005211 = 20000 counts
005251 = 17000 counts
005337 = 15000 counts
005417 = 13000 counts
005457 = 11500 counts
005542 = 10000 counts
005719 = 18000 counts
005816 = 15000 counts
005913 = 13000 counts
010008 = 11500 counts
010104 = 9000 counts

V-Bessel on Blue

WILL FILL IN TOMORROW -- RAN OUT OF TIME

01:25 -Attempted dofpia but still a little bright. Zeroed out the optics and will wait for a bit

01:40 Closer to Sunset -- running dofpia using WF10_321. Still getting message: Too many iterations, unable to compute sky. Trying a different field

01:45 Slewing to ACT_0221; running dofpia with much more success. May have been that the reference star was so bright it was throwing off the bg? Or too crowded a field? Not sure.

02:00 lbcrangebal;

02:20 Slew to target "Decprep1". Checking for guiding, tech chip focus corrections, and dithering. Running dofpia first

02:25 Doug calls: GCS switched back to version 2.7.3b1

02:47 Images are looking elongated especially in blue. Using 180 seconds in V-Bessel on Blue and R-Bessel on Red.

See
024831 in blue and 024823 in red

02:55 Re-running dofpia to see if it will help at all. Temp of mirrors = 3.8C; outside 4C

03:00 Olga skypes in from home...

03:25 031102 in B and 031058 in R. Still show quite a bit of elongation

03:30 Trying a different elevation since this is a little far over - slewing to Decprep2.

03:33 Some questions about z11...calling Doug and Chris

03:40 Issue resolved. optics cleared. Running dofpia on this higher elevation field to see if we see elongation

03:50 Running new target -- still see considerable elongation but better than at lower elevation

See: 035249 in blue and 035248 in red

04:15 Olga in control room...

04:25 Running superfoc sequence. A little soft at 1.2" but it's been considerably better

superfoc7.ob
lbcr.20121202.042010.fits
lbcr.20121202.042051.fits
lbcr.20121202.042131.fits
lbcr.20121202.042211.fits
lbcr.20121202.042251.fits
lbcr.20121202.042331.fits
lbcr.20121202.042410.fits

superfoc7_pa180
lbcr.20121202.042809.fits
lbcr.20121202.042849.fits
lbcr.20121202.042929.fits
lbcr.20121202.043009.fits
lbcr.20121202.043049.fits
lbcr.20121202.043129.fits
lbcr.20121202.043211.fits

superfoc7.ob
lbcb.20121202.042007.fits
lbcb.20121202.042057.fits
lbcb.20121202.042146.fits
lbcb.20121202.042235.fits
lbcb.20121202.042324.fits
lbcb.20121202.042413.fits
lbcb.20121202.042503.fits

superfoc7_pa180
lbcb.20121202.042813.fits
lbcb.20121202.042903.fits
lbcb.20121202.042953.fits
lbcb.20121202.043043.fits
lbcb.20121202.043132.fits
lbcb.20121202.043221.fits
lbcb.20121202.043310.fits

Pause for dofpia

superfoc7_pa90
lbcr.20121202.044354.fits
lbcr.20121202.044434.fits
lbcr.20121202.044514.fits
lbcr.20121202.044553.fits
lbcr.20121202.044633.fits
lbcr.20121202.044714.fits
lbcr.20121202.044754.fits

superfoc7_pa270
lbcr.20121202.045102.fits
lbcr.20121202.045141.fits
lbcr.20121202.045222.fits
lbcr.20121202.045303.fits
lbcr.20121202.045343.fits
lbcr.20121202.045423.fits
lbcr.20121202.045503.fits

superfoc7_pa90
lbcb.20121202.044405.fits
lbcb.20121202.044454.fits
lbcb.20121202.044542.fits
lbcb.20121202.044632.fits
lbcb.20121202.044720.fits
lbcb.20121202.044809.fits
lbcb.20121202.044859.fits

superfoc7_pa270
lbcb.20121202.045105.fits
lbcb.20121202.045155.fits
lbcb.20121202.045244.fits
lbcb.20121202.045332.fits
lbcb.20121202.045421.fits
lbcb.20121202.045511.fits
lbcb.20121202.045600.fits

05:00 dofpia

05:05 running RB_rvexin4 at different PAs

05:30 observing Decprep2 again to see if the elongation has improved. Seeing now seems quite good. DOFPIA is reporting sub arcsecond.

05:45 Seeing a lot of coma (z7), clearing optics again, this time we notice a lot of astigmatism when we restart dofpia. We've decided to try a different field that might have a different rotator angle to see if that has any impact on it.

06:00 slewing to Decprep4

061234 LBCB and 061229 LBCR look fantastic. 0.72" on Red; 0.76" on Blue

MODS1

06:30 Change to MODS. Do modsWake, sieveSnap and imSnap. All look good

07:00 Really struggling to collimate here. LOTS of coma; some issues....David calling Doug but end up solving the problem by applying large global offsets

07:45 Finally collimated. Slewing to recheck MODS collimation issue. We see that the starts in blue look a little up/down elongated, but still much improved from the reported IT.

08:04 Taking spectrophotometric data in all configurations. Star is gd71 (labeled with cirrus label since it is not photometric). We are using this to test out the configuration. I haven't seen some of these and 14 and 15 in red and 12 and 13 in blue look a little different to me.

08:45 Slewing to Dennis' target; GCS Seeing is at 1.2 at El = 42; I am still seeing some disturbing elongation here.

09:05 Slewing to another of Dennis' target. Again some disturbing elongation here. Will document in InformationTechnology once I have even minimal downtime, which I don't right now

LUCI1

09:15 Change to LUCI --

09:35 GCS problem -- see nightlog

09:45 Taking a pupil image, Dave calls it ok. Not sure why black dot is not centered though... luci.20121202.0001.fits

10:00 GCS problem resolved - see nightlog

10:02 Removing pupil imager causes camera failure. Dave has ENG gui open and fixes.

10:15 Trying out acq script (M0647.acq) to see if Kai's rotcenter fix will make MOS acq more robust. Unfortunately we see a lot of elongation in the images. Dave reint the detector block but didn't see change. Images are 0002-0004.

10:30 Now we're attempting a sievemask image to see if there is a problem with the optics. Sieve mask images are 0005-0011.

11:00 Finished executing. Used Rotator positions of 408, 228, 138, 318

11:25 Trying acquisition again. Still elongated but we have some data to try to figure out why now.

11:40 Obviously not a great acquisition. The offsets moved the stars but farther from centered in the boxes. Dave provided an offset

12:00 Tried M011zCgr19.acq. Same as above..MOS acquisition with a poor offset. Again Dave's offset worked fairly well.

12:05 Slewing to a different target to try another MOS acquisition

12:35 M009zCgr32.acq. Ran into twilight; data will need to be checked; bg may have been changing too fast for adequate subtraction

1:00 Closing

01:30 Taking masksnap.pro for MODS. Shutting down LBC. Closing up LUCI. Putting MODS to bed for night

-- MichelleEdwards - 01 Dec 2012
Topic revision: r9 - 25 Jul 2014, DouglasFisher
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback